How to calculate the horizontal and vertical components of a 350 air Force pushing downwards and an angle of 55 degrees to the vertical


1
Expert's answer
2022-01-19T06:00:57-0500

"F_h=F\sin\alpha=287~N,"

"F_v=F\cos\alpha=201~N."